# Scaffold-BPE: Enhancing Byte Pair Encoding with Simple and Effective Scaffold Token Removal

### v1

@@ 3.2 Encoding Process
The encoding process of the original BPE encodes a text $T$ into a token representation (i.e., $R$) using the vocabulary $V$ generated by BPE training. Firstly, $R$ is a sequence of smallest unit tokens (i.e., character/byte tokens), obtained by splitting $T$. And then, following the order of tokens in $V$ as merging priority (i.e., tokens added earlier have higher frequency and thus are assigned higher priority to be merged into), token pairs in $R$ are iteratively merged to build the final representation.
Algorithm 2 Scaffold-BPE Encoding Algorithm
A Text $T$, Expanded Vocabulary $E$
Split $T$ into a character/byte token representation (denoted as $R$)
while True do
     /**************** Scaffolding Begins ****************/
     Identify all possible merges $M$ using $E$, ignoring token types
     /***************** Scaffolding Ends *****************/
     if $M$ is empty then
         break
     end if
     Select $m$ which is ranked before the others in $E$ from $M$
     apply $m$ to $R$
end while
/***************** Demolishing Begins *****************/
Demolish all scaffold tokens in $R$ into its shortest non-scaffold child token sequence
/****************** Demolishing Ends ******************/
return $R$
Similarly, the modifications of Scaffold-BPE in the encoding process are very simple. Compared to the original BPE, the expanded vocabulary $E$ is utilized. Both normal tokens and scaffold tokens are merged according to their rank in $E$. Consequently, during the encoding process, the count of different tokens used actually exceeds the predefined vocabulary size (i.e., $N$). And scaffold tokens are employed as intermediate tokens to merge into longer tokens. We term that mechanism as Scaffolding, as shown in Algorithm 2.
Figure 4: The encoding process (i.e., Scaffolding and Demolishing) for the word “zona" and “Arizona".
When no more token pairs can be merged in $R$, the original BPE returns $R$ as the final result. However, due to the introduction of the Scaffolding mechanism in Scaffold-BPE, $R$ may contain scaffold tokens from $S$, potentially increasing the variety of tokens beyond the predefined vocabulary size and exceeding the range of word embeddings that the model can map. To address it, Scaffold-BPE adds one additional step termed as Demolishing at the end of the encoding process. Scaffold-BPE demolishes all scaffold tokens in $R$ into their shortest non-scaffold child token sequences, ensuring that $R$ only consists of tokens from $V$. For example, as shown in Figure 4, the remaining “zona” in $R$ is demolished into “zon” and “a”. After the Demolishing step, Scaffold-BPE returns the final token sequence representation (i.e., $R$) for $T$. Since the shortest non-scaffold child token sequences for all scaffold tokens can be precomputed and stored during the training process, the time complexity of demolishing one token is $O(1)$, making its impact on encoding efficiency negligible.
@@ 4.1 Experimental Setup
Datasets. Our models are trained on the Pile [15] dataset, an 825.18 GiB English text dataset designed for training large scale language models. The Pile is composed of 22 diverse and high-quality datasets, the models trained on which significantly outperform both raw and filtered Common Crawl [10] models. The data distribution for our model training is identical to those described in the original work [15].
Tokenizer. We train two 32K vocabularies (size applied by LLaMA series [42, 43]) using the original BPE and Scaffold-BPE, respectively. The training text corpus is sampled from the Pile dataset with an identical data distribution. Following GPT-2 [33], we pre-segment the text using its regular expression.
Table 1: Model sizes and architectures.
@@
Model. We train three generative language models with 468M, 1.2B, and 6.7B parameters, respectively. Specifically, the architectures of the 468M-parameter and the 1.2B-parameter models, including the dimensionality of hidden states, the number of layers, etc., are identical to those of the 410M-parameter and the 1.0B-parameter models outlined in Pythia [5]. The minor differences in parameter sizes are attributed to the variations in vocabulary size in the embedding layer. As for the 6.7B-parameter model, its architecture is identical to LLaMA-7B [42]. The corresponding hyperparameters for each model can be found in Table 1.
Training. Following LLaMA [42], we use the AdamW optimizer [28] with a learning rate of $3.0\times 10^{-4}$, $2k$ warmup steps, and a cosine learning rate decay schedule. Following the pretraining settings of previous works [48, 41] and limited by our computation budget, by default all models are pretrained with 100B tokens. Note that the volume of corresponding text data contained in an equal amount of tokens is slightly different between the two tokenizers. Considering model training efficiency and commonly used criteria (i.e., the token amount) of computation budget in LLM training, we still compare experiments in the setting of an equal amount of training tokens. In Section 5.3, we further analyze both tokenizers in the setting of an equal amount of training text volume.
@@
Evaluation. For fair comparisons, we utilize the open-source pipeline lm-evaluation-harness [16] for evaluation [5, 41].
4.2 Experimental Results
Table 2: At varying model scales, the zero-shot and few-shot performance (i.e., accuracy %) on common sense reasoning benchmarks.
HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
468M
0-shot Original BPE 40.81 30.20 66.74 41.62 63.70 53.12
Scaffold-BPE 41.23 30.64 66.94 42.14 63.56 53.20
5-shot Original BPE 40.99 29.20 66.60 43.88 63.00 53.40
Scaffold-BPE 41.19 29.88 66.79 44.98 63.52 53.48
1.2B
0-shot Original BPE 47.21 30.68 69.25 41.77 65.92 55.60
Scaffold-BPE 47.50 31.12 69.52 41.97 66.31 55.53
5-shot Original BPE 47.39 32.36 69.61 46.03 65.79 55.33
Scaffold-BPE 47.64 32.40 69.84 46.58 66.57 56.15
6.7B
0-shot Original BPE 60.84 36.56 73.72 43.00 71.20 61.50
Scaffold-BPE 60.50 37.12 73.99 43.91 71.61 61.97
5-shot Original BPE 61.24 36.24 74.40 49.58 72.63 62.83
Scaffold-BPE 61.34 38.08 74.44 49.89 72.76 63.47
Common Sense Reasoning. Our analysis incorporates six benchmark datasets recognized for evaluating common sense reasoning including HellaSwag [51], OpenBookQA [29], PIQA [6], SIQA [37], StoryCloze [30], and Winogrande [35]. We present the performance of our model, focusing on accuracy in both zero-shot and few-shot scenarios.
As shown in Table 2, we can observe that the Scaffold-BPE consistently outperforms the original BPE on different setups with different model sizes. Specifically, for all models on 5-shot settings, Scaffold-BPE consistently achieves superior performance compared to the original BPE. As for the 0-shot settings, Scaffold-BPE still yields higher performance on 5 out of the 6 datasets.
Table 3: Zero-shot and few-shot exact match performance of 6.7B-parameter language models on closed-book question-answering benchmarks.
TriviaQA WebQuestions
0-shot Original BPE 11.32 3.84
Scaffold-BPE 12.90 5.11
5-shot Original BPE 23.50 16.44
Scaffold-BPE 23.68 17.13
Closed Book Question Answering. For the task of closed book question answering [8, 42, 41], we evaluate the performance of the largest 6.7B-parameter models with different tokenizers on two benchmark datasets, i.e., TriviaQA [22] and WebQuestions [4]. We report the exact match performance for the zero-shot and few-shot settings in Table 3. It can be seen that language models trained with the proposed Scaffold-BPE achieve superior performance in both settings, which demonstrates that Scaffold-BPE can enhance model performance across different types of downstream tasks.
5 Discussion
5.1 Robustness In Language Modeling Tasks
Various Vocabulary Size. Depending on the size of the training corpus, the diversity of the languages, the size of the model, and the types of tasks, different vocabulary sizes are set in practice. Therefore, to validate the robustness of Scaffold-BPE across various vocabulary sizes, in addition to the 32K vocabulary [42, 34], we also trained two vocabularies sized at 64K [3, 2] and 128K [50]. The experimental setup is identical to that of the 468M-parameter model mentioned in Section 1.
Figure 5: The average frequencies of the scaffold tokens and the new actual high-frequency tokens that replace the removed scaffold tokens in vocabularies of 32K, 64K and 128K.
As shown in Table 4, the experiments demonstrate that Scaffold-BPE consistently outperforms the original BPE across all vocabulary sizes, which indicates that Scaffold-BPE is not sensitive to vocabulary size. Its algorithmic design enables it to adaptively identify and remove scaffold tokens across any vocabulary size, without the need for manually designed or heavily-tuned hyperparameters.
Table 4: At varying vocabulary sizes, the zero-shot and few-shot performance (i.e., accuracy %) on common sense reasoning benchmarks.
HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
64K
0-shot Original BPE 41.61 30.00 67.70 41.46 63.45 52.77
Scaffold-BPE 42.09 31.20 67.71 41.80 63.83 52.80
5-shot Original BPE 42.05 30.08 67.58 44.22 63.98 53.25
Scaffold-BPE 42.64 29.92 67.70 45.16 64.40 53.67
128K
0-shot Original BPE 43.59 31.60 67.97 41.72 65.07 53.43
Scaffold-BPE 43.41 31.92 68.80 42.18 65.23 54.30
5-shot Original BPE 43.40 29.24 68.12 45.04 65.08 53.53
Scaffold-BPE 43.45 31.36 69.16 45.91 65.59 54.90
More Training Tokens. According to the Scaling Law, the loss scales as a power-law with model size, dataset size, and the amount of compute used for training [23]. To demonstrate the effectiveness of our Scaffold-BPE with more training tokens, we continue training the 468M-parameter models up to 300B tokens [52, 5, 34, 8].
As shown in Table 5, the experiments demonstrate that Scaffold-BPE consistently outperforms the original BPE at 300B training tokens. The result well indicates that in the era of increasingly large training datasets for Large Language Models, our Scaffold-BPE can effectively enhance the capabilities of those models through simple modifications to the original BPE.
Table 5: At 300B training tokens, the zero-shot and few-shot performance (i.e., accuracy %) on common sense reasoning benchmarks.
HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
0-shot Original BPE 45.56 30.40 68.35 42.08 65.25 54.40
Scaffold-BPE 45.69 30.60 69.53 42.37 65.36 54.95
5-shot Original BPE 45.72 30.72 68.31 46.10 65.58 54.96
Scaffold-BPE 45.78 31.04 69.53 46.15 65.95 56.13
5.2 Task Insensitive, Language Insensitive And Architecture Insensitive
Although the development of large language models is burgeoning and they are increasingly applied across various scenarios, many applications still prefer using conventional models due to their lower training and inference costs. In the field of Natural Language Processing, BPE was initially applied to machine translation tasks [40], which typically present an open vocabulary challenge and involve substantial textual variation between two languages. Consequently, numerous improvements to BPE have been extensively validated on machine translation tasks [32, 49, 20, 45, 12, 46, 36].
Therefore, to validate the versatility of the Scaffold-BPE method, we additionally conduct evaluations on machine translation tasks. We replicate the experimental setup of the prior work [31] which uses 32K vocabularies for the WMT’14 English-German dataset and 40K vocabularies for the WMT’14 English-French dataset [7]. For fair comparisons, We do not pre-segment the text using regular expressions. We train the “big" transformer models [44, 31] to convergence and average model parameters from the last 10 checkpoints [31].
Table 6: BLEU on newstest2014 for WMT English-German (En–De) and English-French (En–Fr).
En-De En-Fr
Original BPE [31] 29.3 43.2
Scaffold-BPE 29.8 43.8
As shown in Table 6, Scaffold-BPE outperforms the original BPE in machine translation tasks, which demonstrates that Scaffold-BPE is not specific to language modeling tasks and can be applied to a wider range of tasks like language understanding, summarization and text classification.
Besides, experiments conducted with English-German and English-French language pairs demonstrate that Scaffold-BPE is language insensitive. Scaffold-BPE is capable of identifying and removing the scaffold tokens introduced by the original BPE across different languages.
Furthermore, prior experiments on language modeling tasks are carried out on decoder-only architecture. For the machine translation tasks, we utilize the classic encoder-decoder architecture [44]. The exceptional performance of Scaffold-BPE confirms its architecture insensitivity, indicating its applicability across a wider range of neural network architectures.
5.3 Better Text Representation
Higher Entropy, Lower Redundancy. Scaffold-BPE can alleviate the imbalance in token frequency, which can lead to an increase in information entropy. We measure Shannon Entropy and Redundancy [18] over token representations of texts obtained with the original BPE and our Scaffold-BPE. Both take as input a text $T$ with a vocabulary of (normal) tokens $V=\{t_{1},t_{2},...,t_{V}\}$ of size $|V|$.
$H(T)=-\sum_{i=1}^{V}p(t_{i})\log_{2}p(t_{i})$ (4)
$R(T)=1-\frac{H(T)}{max\{H(T)\}}=1-\frac{H(T)}{\log_{2}|V|}$ (5)
As shown in Table 7, taking the 32K vocabulary as an example, our Scaffold-BPE can encode Pile dataset [15] with higher Entropy and lower Redundancy. Consequently, tokens in the vocabulary of our Scaffold-BPE have more balanced appearing probabilities. According to Su et al. [41], our vocabulary with balanced token appearances mitigates the learning imbalance problem, resulting in more sufficient learning towards the text corpus, thus achieving better performance.
Table 7: Entropy and Redundancy on tokenized Pile dataset.
[…]
Original BPE 11.2382 0.2491
Scaffold-BPE 11.2443 0.2487
Table 8: Compression Rate (bytes per token) on the Pile dataset and the WMT dataset.
Pile-32K Pile-64K Pile-128K En-De En-Fr
Original BPE 3.879 4.083 4.228 4.830 5.012
Scaffold-BPE 3.889 4.089 4.232 4.861 5.042
Table 9: In the condition of exactly 388 GiB training text, the zero-shot and few-shot performance (i.e., accuracy %) on common sense reasoning benchmarks.
HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
0-shot Original BPE 40.80 30.21 66.74 41.68 63.65 53.12
Scaffold-BPE 41.24 30.66 66.92 42.15 63.55 53.22
5-shot Original BPE 40.97 29.22 66.63 43.86 63.03 53.44
Scaffold-BPE 41.20 29.87 66.78 45.01 63.54 53.50
Higher Compression Rate. Besides the performance of models on tasks, the compression rate is a metric to measure the effectiveness of a tokenizer. A higher compression rate means that fewer tokens are required to represent the same corpus. As shown in Table 8, Scaffold-BPE, utilizing a dynamic scaffold tokens removal mechanism, retains more actual high-frequency tokens in the final vocabulary. Therefore it can achieve a higher compression rate on the corpus.
Besides, considering model training efficiency and commonly used criteria (i.e., the token amount) of computation budget in LLM training, experiments in Section 4.2 and 5.1 are compared in the setting of an equal amount of training tokens. To eliminate the impact of different amounts of training text caused by different compression rates on experiment results, we additionally train two 468M-parameter models on exactly 388 GiB training text ($\approx$ 100B tokens). As shown in Table 9, the Scaffold-BPE consistently outperforms the original BPE, demonstrating that the effectiveness of Scaffold-BPE is not merely obtained by allowing models to digest more data in the same computation budget. Our Scaffold-BPE addresses the issue of token frequency imbalance, allowing models to learn the information contained in tokens more sufficiently, thus achieving better performance.
Better Uniformity of Learned Embeddings. Prior works have analyzed the embedding space learned by a model [32] and found that better uniformity prefers a token embedding space that preserves maximal information [47]. To demonstrate our Scaffold-BPE can mitigate token frequency distribution imbalance, thus leading to a better learned token embedding space with better uniformity, we visualize the token embeddings in the 6.7B-parameter models, following Provilkov et al. [32]. As shown in Figure 6, the embeddings of scaffold tokens learned via the original BPE are more clustered, which means they are not well learned. On the contrary, the embeddings of new tokens introduced by Scaffold-BPE after removing scaffold tokens have better uniformity, which are more evenly distributed across the semantic space. Therefore, models trained with Scaffold-BPE can achieve better performance.
(a) The Original BPE
(b) The Scaffold-BPE
6 Conclusions
In this paper, we present our observation of tokens with imbalanced frequencies in BPE vocabulary, which we term scaffold tokens. Those scaffold tokens, while integral to the formation of longer tokens, do not represent actual frequent tokens in the corpus and affect NLP task performance. To address that issue, we propose Scaffold-BPE, which can remove scaffold tokens from the final token representations by dynamically marking scaffold tokens in the training process and temporarily utilizing scaffold tokens in the encoding process. The Scaffold-BPE is parameter-free, computation-light, easy-to-implement, and widely effective, well preserving the simplicity and clarity of BPE. Through extensive experiments, including varying model sizes, varying vocabulary sizes and extending training tokens, Scaffold-BPE demonstrates its robustness and superiority over the original BPE across a variety of natural language processing tasks. Our work underscores the importance of continual refinement in tokenization methods for improving the overall efficiency and effectiveness of models in the natural language processing field.
7 Limitations
In the proposed Scaffold-BPE, the modifications to the training and encoding of the original BPE are simple and straightforward. Therefore Scaffold-BPE may be combined with other enhancements such as optimal vocabulary size search and novel encoding methods to achieve further improvements. We leave the investigation to our future research.

### v2

@@ 3.2 Encoding Process
The encoding process of the original BPE encodes a text $T$ into a token representation (i.e., $R$) using the vocabulary $V$ generated by BPE training. Firstly, $R$ is a sequence of smallest unit tokens (i.e., character/byte tokens), obtained by splitting $T$. And then, following the ranks of tokens in $V$ as merging priority (i.e., tokens added earlier have higher frequency and thus are assigned higher priority to be merged into), token pairs in $R$ are iteratively merged to build the final representation.
Similarly, the modifications of Scaffold-BPE in the encoding process are straightforward. Compared to the original BPE, the expanded vocabulary $E$ is utilized. In each iteration, the token $t$ to be merged would be selected from both normal tokens and scaffold tokens:
$t=\arg\min_{t\in E}rank_{E}(t)$ (5)
where $rank_{E}(\cdot)$ denotes the rank of a token in $E$. Consequently, during the encoding process, the count of different tokens used actually exceeds the predefined vocabulary size (i.e., $N$). And scaffold tokens are employed as intermediate tokens to merge into longer tokens. We term such a mechanism as Scaffolding, as shown in Algorithm 2.
When no more token pairs can be merged in $R$, the original BPE returns $R$ as the final result. However, due to the introduction of the Scaffolding mechanism in Scaffold-BPE, $R$ may contain scaffold tokens from $S$, potentially increasing the variety of tokens beyond the predefined vocabulary size and exceeding the range of word embeddings that the model can map. To address it, Scaffold-BPE adds one additional step termed as Demolishing at the end of the encoding process. Scaffold-BPE demolishes all scaffold tokens in $R$ into their shortest non-scaffold child token sequences, ensuring that $R$ only consists of tokens from $V$. For example, as shown in Figure 4, the remaining “zona” in $R$ is demolished into “zon” and “a”. The demolishing step can be formulated as follows:
$t=\begin{cases}t,&\text{if Scaffold}(t)=\text{False}\\ (a,b),&\text{otherwise}\end{cases}$ (6)
where $a$ and $b$ are the components of the scaffold token $t$. The formula above would be recursively applied to $a$ and $b$ to derive the shortest non-scaffold child token sequence for $t$. After the Demolishing step, Scaffold-BPE returns the final token sequence representation (i.e., $R$) for $T$. Since the shortest non-scaffold child token sequences for all scaffold tokens can be precomputed during the training process, the time complexity of demolishing one token is $O(1)$, making its impact on encoding efficiency negligible.
Figure 4: The encoding process (i.e., Scaffolding and Demolishing) for the word “zona” and “Arizona”.
@@ 4.1 Experimental Setup
Our models are trained on the Pile (Gao et al. 2020) dataset, an 825.18 GiB English text dataset designed for training LLMs. The data distribution for our model training is identical to that of the original work (Gao et al. 2020).
Tokenizer.
We train two 32K vocabularies (size applied by LLaMA series (Touvron et al. 2023a, b)) using the original BPE and Scaffold-BPE, respectively. Similar to GPT-2 (Radford et al. 2019), pre-tokenization was employed to prevent the merging of tokens from different character categories. And following (Touvron et al. 2023a), we split numbers into individual digits.
Model.
We train three language models with 468M, 1.2B, and 6.7B parameters, respectively. Specifically, the architectures of the 468M and the 1.2B models are identical to those of the 410M and the 1.0B models outlined in Pythia (Biderman et al. 2023). The minor differences in parameter sizes are attributed to the variations in vocabulary size in the embedding and output layer. As for the 6.7B model, its architecture is identical to LLaMA-7B (Touvron et al. 2023a).
Training.
Following the pretraining settings of previous works (Xie et al. 2023; Su et al. 2024; Xiong et al. 2024) and limited by our computation budget, by default all models are pretrained with 100B tokens. Note that the volume of corresponding text data contained in an equal amount of tokens is slightly different between the two tokenizers. Considering model training efficiency and commonly used criteria (i.e., the token amount) of computation budget in LLM training, we compare experiments in the setting of an equal amount of training tokens. In the Section 4.3, we further analyze both tokenizers in the setting of an equal amount of training text volume.
4.2 Experimental Results
Common Sense Reasoning.
Our analysis incorporates 7 benchmarks recognized for evaluating common sense reasoning, including BoolQ (Clark et al. 2019), HellaSwag (Zellers et al. 2019), OpenBookQA (Mihaylov et al. 2018), PIQA (Bisk et al. 2020), SIQA (Sap et al. 2019), StoryCloze (Mostafazadeh et al. 2016), and Winogrande (Sakaguchi et al. 2021). We present the performance of all models in terms of average accuracy in 0-shot and 5-shot settings.
As shown in Table 1, we can observe that the Scaffold-BPE consistently outperforms the original BPE on different setups with different model sizes. Notably, the 6.7B model trained with Scaffold-BPE can achieve a significant 2.08pp (percent point) improvement on BoolQ and a 2.90pp improvement on OpenBookQA. We conduct a $t$-test, and all metrics have $p$-values less than 0.01, indicating that the results are statistically significant.
TriviaQA WebQuestions
Original BPE 15.63 8.56
Scaffold-BPE 18.86 9.89
Table 2: The average exact match performance on 0/5-shot closed-book question-answering benchmarks of the 6.7B-parameter models ($p$-value $<0.01$).
Figure 5: The average frequencies of the scaffold tokens and the new actual high-frequency tokens that replace the scaffold tokens in vocabularies of 32K, 64K and 128K.
Closed Book Question Answering.
For the task of closed book question answering (Brown et al. 2020; Touvron et al. 2023a), we evaluate the performance of the largest 6.7B-parameter models with different tokenizers on 2 benchmark datasets, i.e., TriviaQA (Joshi et al. 2017) and WebQuestions (Berant et al. 2013). We report the exact match performance for the zero-shot and few-shot settings in Table 2. It can be seen that the model trained with the proposed Scaffold-BPE can achieve a 3.23pp improvement on TriviaQA and a 1.33pp improvement on WebQuestions, with both $p$-values less than 0.01. All results above demonstrate that Scaffold-BPE can enhance model performance across different types of downstream tasks.
4.3 Discussion
Various Vocabulary Size.
Depending on the size of the training corpus, the diversity of the languages, the size of the model, and the types of tasks, different vocabulary sizes are set in practice. Therefore, to validate the robustness of Scaffold-BPE across various vocabulary sizes, in addition to the 32K vocabulary (Touvron et al. 2023a), we also trained two vocabularies sized at 64K (Baichuan 2023b, a) and 128K (Yang et al. 2023). The experimental setup is identical to that of the 468M-parameter model mentioned before.
64K 128K
Orig. BPE Scaffold-BPE Orig. BPE Scaffold-BPE
BoolQ 58.01 59.71 56.67 59.43
HellaSwag 41.82 42.06 42.70 42.91
OpenBookQA 30.90 31.10 31.10 32.30
PIQA 67.95 69.26 67.68 68.82
SIQA 43.47 43.86 43.83 44.14
StoryCloze 64.19 65.02 64.11 65.26
Winogrande 53.67 54.34 53.91 55.09
Table 3: At varying vocabulary sizes, the average accuracy on 0/5-shot common sense reasoning benchmarks ($p$-value $<0.01$).
Original BPE Scaffold-BPE
BoolQ 58.21 61.62
HellaSwag 45.10 46.03
OpenBookQA 31.10 33.50
PIQA 68.63 69.86
SIQA 43.78 44.73
StoryCloze 65.53 66.54
Winogrande 53.67 56.12
Table 4: At 300B training tokens, the average accuracy on 0/5-shot common sense reasoning benchmarks ($p$-value $<0.01$).
Moreover, as shown in Table 3, the results demonstrate that Scaffold-BPE consistently outperforms the original BPE across all vocabulary sizes, which indicates that the superiority of Scaffold-BPE is not sensitive to vocabulary size. Its algorithmic design enables it to adaptively remove scaffold tokens across any vocabulary size, without the need for manually designed or heavily-tuned hyperparameters.
More Training Tokens.
According to the Scaling Law, the loss scales as a power-law with model size, dataset size, and the amount of training computation (Kaplan et al. 2020). To demonstrate the effectiveness of our Scaffold-BPE with more training tokens, we continue training the 468M models up to 300B tokens (Zhang et al. 2022; Biderman et al. 2023).
As shown in Table 4, the results demonstrate that Scaffold-BPE consistently outperforms the original BPE at 300B training tokens, well indicating that in the era of increasingly large training datasets for LLMs, our Scaffold-BPE can effectively enhance the capabilities of those models through simple modifications to the original BPE.
Applicable for Other Tasks, Languages, Model Architectures and Compatible with Other BPE Enhancements.
Although the development of LLMs is burgeoning, some applications still prefer using conventional models due to their lower training and inference costs. In the NLP field, BPE was initially combined with transformer models and applied to machine translation tasks (Sennrich et al. 2015), which typically face an open vocabulary challenge and involve substantial textual variations between two languages. Therefore, to validate the versatility of the Scaffold-BPE method, we additionally conduct evaluations on machine translation tasks with identical experimental setup on WMT’14 En-De and En-Fr dataset in the prior work (Ott et al. 2018).
As shown in Table 5, Scaffold-BPE outperforms the original BPE in machine translation tasks, which demonstrates that Scaffold-BPE is not specific to language modeling tasks and can be applied to a wider range of tasks.
Besides, experiments conducted with En-De and En-Fr language pairs demonstrate that Scaffold-BPE is language insensitive. Scaffold-BPE is capable of identifying and removing the scaffold tokens introduced by the original BPE across different languages.
Moreover, previous experiments on language modeling tasks are carried out on the decoder-only architecture. For the machine translation tasks, we utilize the encoder-decoder architecture (Vaswani et al. 2017). The exceptional performance of Scaffold-BPE confirms its architecture insensitivity, indicating its applicability across a wider range of neural network architectures.
Finally, Scaffold-BPE is orthogonal to and can be combined with existing enhancements to BPE, like BPE-Dropout (Provilkov et al. 2019). As shown in Table 5, Scaffold-BPE with BPE-Dropout achieves further improvements on BLEU, well indicating the compatibility of Scaffold-BPE.
En-De En-Fr
Original BPE 29.31 43.20
    + BPE-Dropout 29.50 43.44
Scaffold-BPE 29.76 43.81
    + BPE-Dropout 29.78 43.83
Table 5: BLEU on WMT’14 En–De and En–Fr.
Higher Compression Rate.
Besides the performance of models on downstream NLP tasks, the compression rate for a given text corpus is a metric to measure the effectiveness of a tokenizer. A higher compression rate means that fewer tokens are required to represent the same corpus. As shown in Table 6, Scaffold-BPE, utilizing a scaffold tokens removal mechanism, retains more actual high-frequency tokens in the final vocabulary, and thus it achieves a higher compression rate on all the corpus in our experiments.
Experiments under Same Corpus Size
As mentioned before, considering model training efficiency and commonly used criteria (i.e., the token amount) of computation budget in LLM training, experiments above are compared in the setting of an equal amount of training tokens. To eliminate the impact of different amounts of training text caused by different compression rates on experiment results, we additionally train two 468M-parameter models on exactly 388 GiB training text ($\approx$ 100B tokens). As shown in Table 7, Scaffold-BPE consistently outperforms the original BPE, demonstrating that the effectiveness of Scaffold-BPE is not merely obtained by allowing models to digest more data in the same computation budget. Our Scaffold-BPE also alleviates the issue of token frequency imbalance, allowing models to learn all tokens more sufficiently and evenly, thus achieving better performance.
Pile En-De En-Fr
Original BPE 3.879 4.830 5.012
Scaffold-BPE 3.889 4.861 5.042
Table 6: Compression Rate (the average number of bytes per token) on the Pile dataset and the WMT dataset.
Original BPE Scaffold-BPE
BoolQ 58.72 60.55
HellaSwag 40.84 41.69
OpenBookQA 30.55 32.22
PIQA 66.58 68.78
SIQA 43.40 44.13
StoryCloze 62.85 63.08
Winogrande 53.07 54.25
Table 7: At exactly 388 GiB training text, the average accuracy on 0/5-shot common sense reasoning benchmarks ($p$-value $<0.01$).
5 Conclusions
In this paper, we present our observation of tokens with imbalanced frequencies in BPE vocabulary, which we term scaffold tokens. Those scaffold tokens, while integral to the formation of longer tokens, do not represent actual frequent tokens and affect the performance of LLMs negatively. To address that, we propose Scaffold-BPE, which can remove scaffold tokens from the final token representations by dynamically marking scaffold tokens in the training process and temporarily utilizing them in the encoding process. The Scaffold-BPE is parameter-free, computation-light, easy-to-implement, and widely effective, well preserving the simplicity and clarity of BPE. Through extensive experiments, including varying model sizes, vocabulary sizes and more training tokens, etc., Scaffold-BPE demonstrates its robustness and superiority over the original BPE.
Appendix A Additional Discussion
A.1 Higher Entropy, Lower Redundancy
Scaffold-BPE can alleviate the imbalance in token frequency, which can lead to an increase in information entropy. We measure Shannon Entropy and Redundancy (Gutierrez-Vasques et al. 2021) over token representations of texts obtained with the original BPE and our Scaffold-BPE. Both take as input a text $T$ with a vocabulary of (normal) tokens $V=\{t_{1},t_{2},...,t_{V}\}$ of size $|V|$.
(a) The Original BPE
(b) The Scaffold-BPE
$H(T)=-\sum_{i=1}^{V}p(t_{i})\log_{2}p(t_{i})$ (7)
$R(T)=1-\frac{H(T)}{max\{H(T)\}}=1-\frac{H(T)}{\log_{2}|V|}$ (8)
As shown in Table 8, taking the 32K vocabulary as an example, our Scaffold-BPE can encode Pile dataset (Gao et al. 2020) with higher Entropy and lower Redundancy. Consequently, tokens in the vocabulary of our Scaffold-BPE have more balanced appearing probabilities. According to Su et al. (2023), our vocabulary with balanced token occurrences mitigates the learning imbalance problem, resulting in more sufficient learning towards the text corpus, thus achieving better performance.
A.2 Better Uniformity of Learned Embeddings
Prior works have analyzed the embedding space learned by a model (Provilkov et al. 2019) and found that better uniformity prefers a token embedding space that preserves maximal information (Wang and Isola 2020). To demonstrate our Scaffold-BPE can mitigate token frequency distribution imbalance, thus leading to a better-learned token embedding space with better uniformity, we visualize the token embeddings in the 6.7B-parameter models, following Provilkov et al. (2019).
As shown in Figure 6, the embeddings of scaffold tokens learned via the original BPE are more clustered, which means they are not well learned. On the contrary, the embeddings of new tokens introduced by Scaffold-BPE after removing scaffold tokens have better uniformity, which are more evenly distributed across the semantic space. Therefore, models trained with Scaffold-BPE can achieve better performance.
[…]
Original BPE 11.2382 0.2491
Scaffold-BPE 11.2443 0.2487
Table 8: Entropy and Redundancy on tokenized Pile dataset.
@@ A.2 Better Uniformity of Learned Embeddings
Table 9: Model sizes and architectures.
Appendix B Further Experiment Details
B.1 Model Training Hyper-Parameters
We train three language models with 468M, 1.2B, and 6.7B parameters, respectively. Specifically, the architectures are listed in Table 9. Following LLaMA (Touvron et al. 2023a), we use the AdamW optimizer (Loshchilov and Hutter 2017) with a learning rate of $3.0\times 10^{-4}$, $2k$ warmup steps, and a cosine learning rate decay schedule.
B.2 Evaluation
@@
For fair comparisons, we utilize the open-source pipeline lm-evaluation-harness (Gao et al. 2023) for evaluation. We reported the average evaluation results of the last five checkpoints.
B.3 Computing Infrastructure
The experiments were conducted using a cluster of 16 servers, each equipped with 8 NVIDIA H800 GPUs. The total memory available on each server is 600 GB, and each server is powered by a 70-core CPU. The operating system running on these servers is Ubuntu 20.04.6 LTS. We report the version numbers of used packages in Table 10.
Package Version
absl-py 1.4.0
accelerate 0.21.0
datasets 2.14.3
deepspeed 0.10.0
ninja 1.11.1
protobuf 4.23.4
pytorch-triton 2.1.0+e6216047b8
scikit-learn 1.3.0
scipy 1.11.1
torch 2.1.0.dev20230807+cu121
torchaudio 2.1.0.dev20230807+cu121
torchinfo 1.8.0
torchvision 0.16.0.dev20230807+cu121
triton 2.0.0
wandb 0.15.7
flash-attn 2.0.4
zstandard 0.21.0
seaborn 0.12.2
gradio 3.39.0
Table 10: Versions of used packages.